# Burer-Monteiro factorizability of nuclear norm regularized optimization

### v1

@@ Appendix A Pseudo-stationarity
    $r^{*}>0$, $r\geq r^{*}+\lfloor q\rfloor$, and $W^{*}<0$.
Otherwise, such a feasible solution exists.
Proof.
The whole proof is similar to B.4, but since the optimization steps are different, we here present all the details. First, we do the variable change $(x,g,y,v,\tau,w)\leftarrow(x/\lambda,g/\lambda,y/\lambda,v/\lambda,\tau,w/\lambda)$ for C.2, and then get the next optimization problem:
${\begin{array}[]{cl}\!\!\!\!\sup_{\begin{subarray}{c}x,g,y,v\in\mathbb{R}^{m}\\ \tau\in\mathfrak{P}_{m},w\in\mathbb{R}\end{subarray}}&\!\!\!\!\!\!\!\lambda^{2}\bigg{(}\sum_{i=1}^{m}(Lx_{i}+g_{i})v_{\tau(i)}\!+\!\sum_{i=1}^{m}g_{i}(Ly_{\tau(i)}+v_{\tau(i)})\!-\!\sum_{i=1}^{m}(Lx_{i}+g_{i})g_{i}\!-\!\sum_{i=1}^{m}(Ly_{i}+v_{i})v_{i}\bigg{)}\\ \!\!\!\!s.t.&\!\!\!\!\!\!\!\forall i\in[r],~{}x_{i}\geq w>0,~{}g_{i}=1,\quad\forall i\in[r^{*}],~{}0<y_{i}\leq\hat{M},~{}v_{i}=1,~{}\sum_{j=r^{*}+1}^{m}v_{j}\leq q,\\ &\!\!\!\!\!\!\!\forall i\in[m]\setminus[r],~{}x_{i}=0,~{}g_{i}\in[0,1+Lw],\quad\forall i\in[m]\setminus[r^{*}],~{}y_{i}=0,~{}v_{i}\in[0,1],\end{array}}$ (C.27)
where $\hat{M}=\frac{M}{\lambda}$. Next, define the following function $H:\mathbb{N}_{0}\times\mathbb{R}_{+}\to\mathbb{R}$:
$H(d,w):=\begin{cases}(q-d)(Lw+1-\frac{q}{d})+H_{0}(d,w)&{\rm if}\ q<d,\\ H_{0}(d,w)&{\rm otherwise}.\end{cases}$ (C.28)
where $H_{0}(d,w):=(r^{*}-r+d)L^{2}(\hat{M}-\min\{w,\hat{M}/2\})\min\{w,\hat{M}/2\}$. Then in view of (B.14) and (C.4), the optimal value of (C.3) is given by $H(|\mathcal{J}^{\tau}_{2}|,w)$.
Using Lemma C.3 and the definition of $H$ in (C.28), upon dropping $\lambda^{2}$ in the objective function since it would not affect the sign of the function value, we can consider C.29 instead because our claim only concerns the feasible set of C.2 and the sign of its objective value:
$\sup_{w\in\mathbb{R},~{}d\in\mathbb{N}_{0}}H(d,w)\quad\quad s.t.~{}w>0,~{}r-r^{*}\leq d\leq\min\{r,m-r^{*}\},$ (C.29)
where the constraint on $d$ comes from the requirement that $|\mathcal{J}^{\tau}_{i}|\geq 0$ for all $i\in[4]$ (see B.14).
We consider the following scenarios:
  • (S1)
    The optimal value of (C.29) is nonpositive, and (C.29) has no feasible solution $(d,w)$ satisfying $H(d,w)\geq 0$ and $r^{*}-r+d>0$.
    In this scenario, we claim that (C.2) has no feasible solution $(\bar{x},\bar{g},\bar{y},\bar{v},\bar{\tau},\bar{w})$ with $\|\bar{g}\|_{\infty}>\lambda$ and $G(\bar{x},\bar{g},\bar{y},\bar{v},\bar{\tau},\bar{w})\geq 0$. A short proof is provided below.
    Suppose such a feasible solution $(\bar{x},\bar{g},\bar{y},\bar{v},\bar{\tau},\bar{w})$ of (C.2) exists, then either $(\bar{x},\bar{g},\bar{y},\bar{v})/\lambda$ is optimal for (C.3) with $\tau=\bar{\tau}$ and $w=\bar{w}/\lambda$, or $(\bar{x},\bar{g},\bar{y},\bar{v})/\lambda$ is not optimal. In the latter case, the optimal value of (C.29) must be positive. In the former case, we see from Lemma C.3 that $|\mathcal{J}_{3}^{\tau}|>0$, and hence (C.29) has a feasible solution $(\tilde{d},\tilde{w}):=(|\mathcal{J}^{\bar{\tau}}_{2}|,\bar{w}/\lambda)$ with $H(\tilde{d},\tilde{w})\geq 0$ and $r^{*}-r+\tilde{d}=r^{*}-r+|\mathcal{J}^{\bar{\tau}}_{2}|=|\mathcal{J}^{\bar{\tau}}_{3}|>0$ (see (B.14)). Both cases yield a contradiction.
  • (S2)
    There exists a feasible solution $(\tilde{d},\tilde{w})$ of C.29 satisfying $H(\tilde{d},\tilde{w})\geq 0$ and $r^{*}-r+\tilde{d}>0$. In this scenario, (C.2) has a feasible solution $(\bar{x},\bar{g},\bar{y},\bar{v},\bar{\tau},\bar{w})$ with $G(\bar{x},\bar{g},\bar{y},\bar{v},\bar{\tau},\bar{w})\geq 0$ and $\|\bar{g}\|_{\infty}>\lambda$. Indeed, we just need to take a $\bar{\tau}\in\mathfrak{P}_{m}$ with $|\mathcal{J}_{2}^{\bar{\tau}}|=\tilde{d}$, and then take $(\tilde{x},\tilde{g},\tilde{y},\tilde{v})$ to be the optimal solution of (C.3) with $\tau=\bar{\tau}$ and $w=\tilde{w}$, and set $(\bar{x},\bar{g},\bar{y},\bar{v},\bar{w})=\lambda(\tilde{x},\tilde{g},\tilde{y},\tilde{v},\tilde{w})$.
Consider the following cases:
  • Case 1:
    $r^{*}=0$. In this case, we see that every feasible solution $(d,w)$ of (C.29) satisfies $d=r$ and hence $r^{*}-r+d=0$, and the optimal value of (C.29) is nonpositive. Then (S1) holds.
  • Case 2:
    $r^{*}>0$ and $r\leq r^{*}+\lfloor q\rfloor-1$. In this case, we can set $d=r-r^{*}+1$ and $w=\hat{M}/2$, which is feasible due to that $r-r^{*}+1\leq r$ and $r-r^{*}+1\leq\lfloor q\rfloor\leq m-r^{*}$. Then $r^{*}-r+d=1$ and $d\leq\lfloor q\rfloor\leq q$. Thus,
    $H(d,w)=H_{0}(d,w)=(r^{*}-r+d){L^{2}(\hat{M}-\min\{w,\hat{M}/2\})\min\{w,\hat{M}/2\}}=\frac{L^{2}\hat{M}^{2}}{4}.$
    Hence (S2) holds.
  • Case 3:
    $r^{*}>0$ and $r\geq r^{*}+\lfloor q\rfloor$. Consider the following problem:
    ${\begin{array}[]{cl}\!\!\!\sup_{w\in\mathbb{R},d\in\mathbb{N}_{0}}&\!\!\!\!\!(q-d)(1\!+\!wL\!-\!\frac{q}{d})\!+\!L^{2}(r^{*}\!-\!r\!+\!d)(\hat{M}\!-\!\min\{w,\hat{M}/2\})\min\{w,\hat{M}/2\}\\ \!\!\!s.t.&\!\!\!\!\!r-r^{*}+1\leq d\leq\min\{r,m-r^{*}\},~{}w>0;\end{array}}$ (C.30)
    notice that $d\geq r-r^{*}+1\geq\lfloor q\rfloor+1>q$ for all $(w,d)$ feasible for (C.30), and hence the objective of C.30 equals the $H(d,w)$ defined in C.28. Observe that (C.30) is obtained by excluding $d=r-r^{*}$ from (C.29), and we have by direct computation that $H(r-r^{*},w)\leq H_{0}(r-r^{*},w)=0$. Thus, if (C.30) has a feasible solution $(d,w)$ with $H(d,w)\geq 0$, then (S2) holds; otherwise (S1) holds.
    Note the objective in C.30 is nonincreasing on $[\frac{\hat{M}}{2},\infty)$ as a function of $w$ when $d$ is fixed. Hence, it holds that C.30 has a feasible solution $(d,w)$ with $H(d,w)\geq 0$ if and only if the following optimization problem has a feasible solution $(d,w)$ with nonnegative objective value:
    ${\begin{array}[]{cl}\sup_{w\in\mathbb{R},d\in\mathbb{N}_{0}}&H_{1}(d,w):=(q-d)(1+wL-\frac{q}{d})+L^{2}(r^{*}-r+d)(\hat{M}w-w^{2})\\ s.t.&r-r^{*}+1\leq d\leq\min\{r,m-r^{*}\},~{}w\in(0,\hat{M}/2].\end{array}}$ (C.31)
    Let $\alpha=\frac{q-d+L(r^{*}-r+d)\hat{M}}{2L(r^{*}-r+d)}$, and we note that $\alpha<\hat{M}/2$ since $d>q$. We rewrite $H_{1}$ as follows:
    $H_{1}(d,w)$ $=(q-d)(1-\frac{q}{d})+L(L(r^{*}-r+d)\hat{M}+(q-d))w-L^{2}(r^{*}-r+d)w^{2}$
    $=(q-d)(1-\frac{q}{d})-L^{2}(r^{*}-r+d)[(w-\alpha)^{2}-\alpha^{2}].$
    We can now rewrite (C.31) as:
    ${\begin{array}[]{cl}\sup_{w\in\mathbb{R},d\in\mathbb{N}_{0}}&H_{1}(d,w)=(q-d)(1-\frac{q}{d})-L^{2}(r^{*}-r+d)[(w-\alpha)^{2}-\alpha^{2}]\\ s.t.&r-r^{*}+1\leq d\leq\min\{r,m-r^{*}\},~{}w\in(0,\hat{M}/2],\\ &\alpha=\frac{q-d+L(r^{*}-r+d)\hat{M}}{2L(r^{*}-r+d)}.\end{array}}$ (C.32)
    If the choice of $d$ makes $\alpha\leq 0$, then we have $(w-\alpha)^{2}-\alpha^{2}=w^{2}-2\alpha w>0$, which together with $r^{*}-r+d\geq 1>0$ implies that
    $H_{1}(d,w)<(q-d)(1-\frac{q}{d})\overset{\rm(a)}{<}0,$
    where in (a) we have used $d>q$. Therefore, (C.32) has a feasible solution with nonnegative function value if and only if it does so when $\alpha>0$. This leads us to consider the following problem:
    ${\begin{array}[]{cl}\sup_{w\in\mathbb{R},d\in\mathbb{N}_{0}}&(q-d)(1-\frac{q}{d})-L^{2}(r^{*}-r+d)[(w-\alpha)^{2}-\alpha^{2}]\\ s.t.&r-r^{*}+1\leq d\leq\min\{r,m-r^{*}\},~{}w\in(0,\hat{M}/2],\\ &\alpha=\frac{q-d+L(r^{*}-r+d)\hat{M}}{2L(r^{*}-r+d)}>0.\end{array}}$ (C.33)
    Maximizing with respect to $w$ and noticing that $\alpha<\frac{\hat{M}}{2}$, we see that (C.33) and the following problem must have or do not have a feasible solution with nonnegative function value simultaneously:
    ${\begin{array}[]{cl}\sup_{d\in\mathbb{N}_{0}}&(q-d)(1-\frac{q}{d})+L^{2}(r^{*}-r+d)\alpha^{2}\\ s.t.&r-r^{*}+1\leq d\leq\min\{r,m-r^{*}\},~{}\alpha=\frac{q-d+L(r^{*}-r+d)\hat{M}}{2L(r^{*}-r+d)}>0.\end{array}}$ (C.34)
    Simplifying this problem, we obtain
    ${\begin{array}[]{cl}\sup_{d\in\mathbb{N}_{0}}&(q-d)(1-\frac{q}{d})+\frac{(q-d+L\hat{M}(r^{*}-r+d))^{2}}{4(r^{*}-r+d)}\\ s.t.&r-r^{*}+1\leq d\leq\min\{r,m-r^{*}\},~{}q-d+L\hat{M}(r^{*}-r+d)>0.\end{array}}$ (C.35)
    The problem (C.35) is compactly discrete, and hence optimal value can be always achieved as long as it is not negative infinity. Therefore, (C.35) has a feasible point with nonnegative function value if and only if the optimal value of (C.35) is nonnegative, in which case (S2) holds, and otherwise (S1) holds.

### v2

@@ 5 Generalization to the RIP case
    $r>r^{*}$ and $\kappa=3$.
  • (3)
    $r\geq r^{*}$ and $\kappa<3$.
  • (4)
    $r\geq r^{*}$, $\kappa>3$ and $\alpha>\min\{r,m-r^{*}\}$.
Finally, we notice that $r-r^{*}=\min\{r,m-r^{*}\}-\min\{r^{*},m-r\}$, which can be proved by discussing the two cases $r+r^{*}\leq m$ and $r+r^{*}>m$ separately. Then, we have $\alpha=(\min\{r,m-r^{*}\}-\min\{r^{*},m-r\})/(1-\frac{4}{(\kappa-1)^{2}})$. Thus, when $\kappa>3$, we can rewrite the condition $\alpha>\min\{r,m-r^{*}\}$ as
$\min\{r,m-r^{*}\}>\frac{(\kappa-1)^{2}}{4}\min\{r^{*},m-r\}.$ (A.24)
Moreover, when $\kappa=3$, the condition A.24 is the same as $\min\{r,m-r^{*}\}-\min\{r^{*},m-r\}>0$, i.e., $r>r^{*}$. Furthermore, when $\kappa<3$ and $r^{*}\leq r<m$, we have $\frac{(\kappa-1)^{2}}{4}<1$ and hence we always have
$\min\{r,m-r^{*}\}-\min\{r^{*},m-r\}=r-r^{*}>\left(\frac{(\kappa-1)^{2}}{4}-1\right)\min\{r^{*},m-r\};$
this is because the strict inequality holds trivially when $r>r^{*}$, and if $r=r^{*}$, we have $\min\{r^{*},m-r\}=\min\{r,m-r\}>0$, which also implies the strict inequality above. Hence, we see from the above display that A.24 holds trivially when $\kappa<3$ and $r^{*}\leq r<m$. Consequently, the above four cases now can be recapped as the following two cases:
  • (1)
    $r=m$.
  • (2)
    $r\geq r^{*}$ and $\min\{r,m-r^{*}\}>\frac{(\kappa-1)^{2}}{4}\min\{r^{*},m-r\}$.